OpenStudy (anonymous):

(GIVE A MEDAL AND FAN) PLEASE HELP!!! Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

@nerlineg @DaBest21 help please

OpenStudy (anonymous):

help please guys!!!!! @DaBest21 @nerlineg @mathmath333 @Pinkflower guys!!!

OpenStudy (anonymous):

i really need help on this please!!!!!

OpenStudy (anonymous):

@EmilyF03 do u know what standard form is

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

are you there? @DaBest21

OpenStudy (anonymous):

ok standard form is A*x+B*y=C

OpenStudy (anonymous):

okayy so whats nexts?

OpenStudy (anonymous):

is this flvs

OpenStudy (anonymous):

yes :(

OpenStudy (anonymous):

what lesson

OpenStudy (anonymous):

03.11

OpenStudy (anonymous):

k one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B)f(x)=75x+60 C) To graph you would start at (0,60) and go up with a slope of 75 use intervals of 1 on the x axis and intervals of 75 on the y axis.

OpenStudy (anonymous):

ill do a real quick one sec

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

Let (3,285) and (6,510) reprsent the coordinate points, where x = number of days coat is rented for; and y = price of rental Given these 2 points, we could write the equation of the straight line. We need the slope. Slope = (510-285)/(6-3) = 225/3 = 75 So the slope is 75. So the equation of the line is y - 285 = 75(x - 3) y - 285 = 75x - 225 y = 75x + 60

OpenStudy (anonymous):

THAT its for part B? right? @mathway

OpenStudy (anonymous):

the above isnt standard form

OpenStudy (anonymous):

@mathway it says standard form no variables can be on the right

OpenStudy (anonymous):

−75x+y=60 is a correct standard form equation

OpenStudy (anonymous):

and @mathway plz do not copy and paste answers from previous questions about this

OpenStudy (anonymous):

its in the code of conduct

OpenStudy (anonymous):

@DaBest21 so the correct answer for part B its: −75x+y=60

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@DaBest21 It is best to guide them first before giving the answers! and for your quetsion, I lost the connection, so I wasn't able to type it next to my last comment. :/

OpenStudy (anonymous):

Remember that the standard form is ax+by=c

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

and for part C the correct answer its: To graph you would start at (0,60) and go up with a slope of 75 use intervals of 1 on the x axis and intervals of 75 on the y axis. ???

OpenStudy (anonymous):

1y=y

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops forgot the to put the link. again due to the lost of connection. Are we clear now, NOT @DaBest21 ?

OpenStudy (anonymous):

whoa calm down ok

OpenStudy (anonymous):

potassium.

OpenStudy (anonymous):

and for part A? @DaBest21

OpenStudy (anonymous):

-75x+y=60 but if u want to be thechnical do -75x+1y=60 which changes nothing

OpenStudy (anonymous):

i was directing the end of that not at u @EmilyF03

OpenStudy (anonymous):

Part A: -75x + 1y = 60, Part B: f(x)=75x+60 Part C: To graph you would start at (0,60) and go up with a slope of 75 use intervals of 1 on the x axis and intervals of 75 on the y axis. <SO THATS the complete answers?

OpenStudy (anonymous):

??/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!!! @DaBest21 and @mathway